OpenStudy (anonymous):

GIven that \(f(x) = \sqrt{9 - x^2}\), and that \(g(x) = f(x - 4)\), rewrite g(x) in terms of x. Please tell me how to do it and not just the answer.

OpenStudy (anonymous):

\[g(x)=f(x-4)\]Notice that\[f(x-4)=\sqrt{9-(x-4)^2}\] Follow so far?

OpenStudy (anonymous):